### v1

# Ergodic measures with large entropy have long unstable manifolds for $C^{\infty}$ surface diffeomorphisms

Abstract
We prove that for ergodic measures with large entropy have long unstable manifolds for $C^{\infty}$ surface diffeomorphisms: For any $\alpha>0$, there exists $\beta>0$ and $c>0$ such that for every ergodic measure $\mu$ with metric entropy large than $\alpha$, then the set of all points which have the size of unstable manifold large than $\beta$ has $\mu$-measure large than $c$.
1 Introduction
Pesin theory is one of the cornerstones in smooth ergodic theory. It has been started from Pesin in 1970s [15, 16]. One of the main parts of Pesin theory is, for a smooth diffeomorphism $f$ of a closed compact Riemannian manifold $M$, for an ergodic hyperbolic measure $\mu$, $\mu$-almost every point has stable manifold and unstable manifold. The size of stable manifold or unstable manifold depends on the measure.
Recall the Oseledec theorem [14]: for any ergodic measure $\mu$ of a diffeomorphism $f$, there are finite numbers
$\lambda_{1}(\mu,f)>\lambda_{2}(\mu,f)>\cdots>\lambda_{s}(\mu,f)$
and a $Df$-invariant splitting on a full measure set $\mu$:
$E^{1}\oplus E^{2}\oplus\cdots\oplus E^{t}$
such that $\sum_{i=1}^{t}\operatorname{dim}E^{i}=d$, and for any non-zero vector $v\in E^{i}$
$\lim_{n\to\infty}\frac{1}{n}\log\|Df^{n}v\|=\lambda_{i}(\mu,f).$
When there is no confusion, one denote $\lambda_{i}(\mu,f)$ by $\lambda_{i}(\mu)$ simply.
Pesin [15, 16] has proved that for $C^{1+\alpha}$ diffeomorphism $f$, for an ergodic measure $\mu$, if $\lambda_{1}(\mu)>0$, then $\mu$-almost every point has unstable manifold. Take $i_{0}$ be the maximal integer such that $\lambda_{i_{0}}(\mu)>0$, then for $\mu$-almost every point $x$, one has a Pesin unstable manifold
$W^{u}_{\rm Pes}(x):=\left\{y\in M:~{}\limsup_{n\to\infty}\frac{1}{n}\log d(f^{n}(x),f^{n}(y))<0\right\}$
tangent to $E^{1}(x)\bigoplus E^{2}(x)\bigoplus\cdots\bigoplus E^{i_{0}}(x)$ at $x$. Similar conclusion holds for the stable manifold.
Although one knows the existence of stable/unstable manifold of a hyperbolic measure, the size of stable/unstable manifold depends on the measure. For $C^{\infty}$ surface diffeomorphisms, we will prove that the size can be decided by the metric entropy of $\mu$ uniformly.
We give the meaning of large unstable manifolds. Denote by
$E^{u}(x)=E^{1}(x)\bigoplus E^{2}(x)\bigoplus\cdots\bigoplus E^{i_{0}}(x).$
Given $\beta>0$, define
$~{}L^{u}(\beta)=\{x\in M:~{}\exists W_{x}\subset W^{u}_{\rm Pes}(x),~{}\textrm{s.t.}~{}\exp_{x}^{-1}W_{x}~{}\textrm{is a $C^{1}$ graph of a map~{}}\varphi:~{}E^{u}(x)\to(E^{u}(x))^{\perp},$
$~{}~{}~{}~{}{\rm Lip}(\varphi)\leq 1/3,~{}{\rm Domain}(\varphi)\supset E^{u}(x)(\beta)\}.$
Similarly, take $j_{0}$ be the maximal integer such that $\lambda_{j_{0}}(\mu)>0$ and denote by
$E^{s}(x)=E^{j_{0}}(x)\bigoplus E^{j_{0}+1}(x)\bigoplus\cdots\bigoplus E^{t}(x).$
Given $\beta>0$, define
$~{}L^{s}(\beta)=\{x\in M:~{}\exists W_{x}\subset W^{s}_{\rm Pes}(x),~{}\textrm{s.t.}~{}\exp_{x}^{-1}W_{x}~{}\textrm{is a $C^{1}$ graph of a map~{}}\varphi:~{}E^{s}(x)\to(E^{s}(x))^{\perp},$
$~{}~{}~{}~{}{\rm Lip}(\varphi)\leq 1/3,~{}{\rm Domain}(\varphi)\supset E^{u}(x)(\beta)\}.$
Our main theorem in dimension $2$, i.e., $M$ is a closed surface, states that if the entropy is large enough, then we have both long stable manifold and long unstable manifold.
Theorem A.
Suppose that $M$ is a closed surface and $f:~{}M\to M$ is a $C^{\infty}$ surface diffeomorphism. For any $\alpha>0$, there are $\beta>0$ and $c>0$ such that for any ergodic measure $\mu$, if
$h_{\mu}(f)>\alpha,$
then $\mu(L^{u}(\beta))>c$ and $\mu(L^{s}(\beta))>c$.
Note that D. Burguet has observed more accurate dependence of $\alpha$ and $c$ in the main theorems.
By personal communications with S. Crovisier, Theorem A answers one question of Crovisier with an additional assumption: the entropy is uniformly bounded away from 00. Crovisier has asked that for $C^{\infty}$ surface diffeomorphisms, if the Lyapunov exponents of an ergodic measure $\mu$ are uniformly bounded away from 00, then some $\mu$ typical points have long unstable manifolds. See [8, Section 3, Question 11] for its version for Hénon maps. Note that by Ruelle inequality, for surface diffeomorphisms, if the entropy is bounded away from 00, then the Lyapunov exponents are bounded away from zero.
The above Theorem A has its $C^{r}$ version. Define
$R(f)=\lim_{n\to\infty}\frac{1}{n}\log\|Df^{n}\|_{\rm sup}.$
Theorem B.
Suppose that $M$ is a closed surface. Let $1<r\in{\mathbb{N}}$. $f:~{}M\to M$ is a $C^{r}$ surface diffeomorphism. For any $\alpha>0$, there are $\beta>0$ and $c>0$ such that for any ergodic measure $\mu$, if
$h_{\mu}(f)>\alpha+\frac{\max\{R(f),R(f^{-1})\}}{r},$
then $\mu(L^{u}(\beta))>c$ and $\mu(L^{s}(\beta))>c$.
For some $C^{r}$ dissipative surface diffeomorphism, a condition on the Jacobian of the map can imply long stable manifold, but not long unstable manifold. See [6, 13] for instance.
It is clear that Theorem A follows from Theorem B. These two theorems are possible to get both long stable manifold and long unstable manifold, since in dimension $2$, by applying the Ruelle inequality [18], every ergodic measure $\mu$ with positive entropy is hyperbolic, i.e., $\lambda_{i}(\mu)\neq 0$ for any $1\leq i\leq s$ in the Oseledec theorem.
Note that Buzzi-Crovisier-Sarig [5] proved the following remarkable result: for $C^{\infty}$ surface diffeomorphisms with positive topological entropy, measures of maximal entropy have the strongly positive recurrence property, i.e., there is a Pesin block such that for ergodic measures close to the measures of maximal entropy, the Pesin block has large measures. Consequently, they proved that for ergodic measures close to the measures of maximal entropy, there are points having both long stable manifolds and long unstable manifolds simultaneously. With the “simultenuous” property, they have very nice properties: the ergodic measures are homoclinically related, and can be coded in one topological Markov shift, etc. In our theorem, we can know the existence of long stable manifold and long unstable manifold, but not simultaneously. So we do not have other consequences like theirs. However, in our main theorems, we do not have to assume the measures are close to the measures of maximal entropy and the size of unstable manifold is uniform with respect to the value of the metric entropy.
We also learned that M. Ghézal has proved that for $C^{2}$ diffeomorphisms, when the metric entropy is larger than some dynamical quantity, some typical points have both long stable manifold and long unstable manifold simultaneously.
One of the main tools of the proof of Theorem B is based on Yomdin theory [20], and its recent important progress by Burguet [3, 4]. It mainly uses Yomdin theory for $1$-dimensional curves. Thus, some corresponding versions of Theorem B and Theorem A hold in the higher dimensional case if we deal with the $1$-dimensional entropy.
From the Oseledec theorem stated above, for an ergodic measure $\mu$ with positive Lyapunov exponents, one says that $\mu$ has one-dominated Lyapunov exponent, if $\operatorname{dim}E^{1}=1$ as $E^{1}$ in the Oseledec theorem. In this case, by the Pesin theory, for $C^{2}$ diffeomorphism $f$, one has one-dimensional strong unstable manifold $W^{1}(x)$ for $\mu$-almost every point $x$. To describe the complexity of the dynamics along $W^{1}$, one can define the partial entropy $h^{1}_{\mu}(f)$ as in [11]. In the higher-dimensional case, we can prove that if this partial entropy is large, then the size of $W^{1}$ can be uniformly long. Similar to $L^{u}(\beta)$, one defines $L^{u,1}(\beta)$:
$~{}L^{u,1}(\beta)=\{x\in M:~{}\exists W_{x}\subset W^{1}(x),~{}\textrm{s.t.}~{}\exp_{x}^{-1}W_{x}~{}\textrm{is a $C^{1}$ graph of a map~{}}\varphi:~{}E^{1}(x)\to(E^{1}(x))^{\perp},$
$~{}~{}~{}~{}{\rm Lip}(\varphi)\leq 1/3,~{}{\rm Domain}(\varphi)\supset E^{1}(x)(\beta)\}.$
Theorem C.
$M$ is a compact Riemannian manifold without boundary of any dimension. Let $1<r\in{\mathbb{N}}$ and $f:~{}M\to M$ is a $C^{r}$ diffeomorphism. For any $\alpha>0$, there is $\beta>0$ and $c>0$ such that for any ergodic measure $\mu$ with one dominated Lyapunov exponent, if
$h^{1}_{\mu}(f)>\alpha+\frac{R(f)}{r},$
then $\mu(L^{u,1}(\beta))>c$.
Theorem D.
$M$ is a compact Riemannian manifold without boundary of any dimension. $f:~{}M\to M$ is a $C^{\infty}$ diffeomorphism. For any $\alpha>0$, there is $\beta>0$ and $c>0$ such that for any ergodic measure $\mu$ with one dominated Lyapunov exponent, if
$h^{1}_{\mu}(f)>\alpha,$
then $\mu(L^{u,1}(\beta))>c$.
Clearly Theorem D can be obtained as a corollary of Theorem C. Theorem C can imply Theorem B.
Remark that the measurability of $L^{u}(\beta)$ and $L^{u,1}(\beta)$ will be checked in Appendix A.
Another interesting subject is the flow $(\varphi^{t})_{t\in\mathbb{R}}$ generated by vector field $X$ over a compact Riemannian manifold $M$ without boundary. In contrast to diffeomorphisms, to study dynamics of flows, usually the difficulties come from the existence of singularities, where the vector field vanishes. See for instance [12, 7].
However, if we want to establish similar theorems for vector fields, it seems the singularities do not give any trouble because to consider the time-one map of the flow is sufficient to get a result similar to above theorems. For instance, we have the following theorem for three-dimensional vector fields:
Theorem E.
Suppose that $M$ is a three-dimensional compact Remiannian manifold without boundary. $X:~{}M\to M$ is a $C^{\infty}$ vector field and $(\varphi^{t})_{t\in\mathbb{R}}$ is the flow generated by $X$. For any $\alpha>0$, there are $\beta>0$ and $c>0$ such that for any ergodic measure $\mu$, if
$h_{\mu}(X)=h_{\mu}(\varphi^{1})>\alpha,$
then $\mu(L^{u}(\beta))>c$ and $\mu(L^{s}(\beta))>c$.
Note that Theorem E is a consequence of Theorem D. Thus the main work of this paper is to prove Theorem C.
Acknowledgments
We are grateful to D. Burguet for his lectures and his comments, and to S. Crovisier for his comments.
2 The entropy theory
For a probability measure $\mu$ (not necessarily invariant), for a finite partition $\mathcal{P}$, define the static entropy of $\mu$:
$H_{\mu}(\mathcal{P})=\sum_{P\in\mathcal{P}}-\mu(P)\log\mu(P)=\int-\log\mu(P(x)){\rm d}\mu(x).$
For a diffeomorphism $f$, one defines
${\mathcal{P}}^{n}=\bigvee_{j=0}^{n-1}f^{-j}(\mathcal{P}).$
For an invariant measure $\mu$, the metric entropy of $\mu$ with respect to a partition $\mathcal{P}$ is
$h_{\mu}(f,\mathcal{P})=\lim_{n\to\infty}\frac{1}{n}H_{\mu}({\mathcal{P}}^{n});$
and the metric entropy of $\mu$ is defined to be
$h_{\mu}(f)=\sup_{\mathcal{P}:~{}\textrm{a partition}}h_{\mu}(f,\mathcal{P}).$
Note that Theorem C and Theorem D are results on “entropy along an invariant foliation” or “partial entropies” as in Ledrappier-Young [11].
Recall the Oseledec theorem as in the introduction, consider $i_{0}$ such that $\lambda_{i_{0}}>0$. As noticed by [19, 11], by denoting $E=E^{1}\bigoplus E^{2}\bigoplus\cdots\bigoplus E^{i_{0}}$, define
$W^{E}(x):=\left\{y\in M:~{}\limsup_{n\to\infty}\frac{1}{n}\log d(f^{-n}(x),f^{-n}(y))\leq-\lambda_{i_{0}}\right\}.$
Then $W^{E}(x)$ is a $C^{2}$ $\operatorname{dim}E$-dimensional immersed submanifold of $M$ tangent to $E(x)$ at $x$ for $\mu$-almost every point $x$. Each $W^{E}(x)$ inherits a Riemannian metric from $M$. The distances are denoted by $d^{E}_{x}$. With this distance, one can define $(n,\rho)$-Bowen balls:
$V^{E}(x,n,\rho):=\{y\in W^{E}(x):~{}d^{E}_{f^{j}(x)}(f^{j}(x),f^{j}(y))<\rho,~{}\forall 0\leq j<n\}.$
From [10], one knows there exists a measurable partition $\xi$ subordinate to $W^{E}$, i.e., for almost every point $x$, one has that $\xi(x)\subset W^{E}(x)$ and contains an open neighborhood of $x$ in $W^{E}(x)$.
Then from Rokhlin [17], associated to each measurable partition $\xi$, there is a system of conditional measures $\{\mu_{\xi(x)}\}$.
As in [11], the partial entropy along $W^{E}$ with respect to $\xi$ is defined to be
$h^{E}_{\mu}(x,\xi,f)=\lim_{\rho\to 0}\liminf_{n\to\infty}-\frac{1}{n}\log(\mu_{\xi(x)}(V^{E}(x,n,\rho)))=\lim_{\rho\to 0}\limsup_{n\to\infty}-\frac{1}{n}\log(\mu_{\xi(x)}(V^{E}(x,n,\rho))).$ (1)
The limits exist as proved in [11]. Also noticed in [11], for an ergodic measure $\mu$, $h^{E}_{\mu}(x,\xi,f)$ does not depend on $x$; and furthermore, it does not depend on the choice of the measurable partition $\xi$. Hence the partial entropy along $W^{E}$ for an ergodic measure $\mu$ is defined to be $h^{E}_{\mu}(x,\xi,f)$, and is denoted by $h^{E}_{\mu}(f)$.
Also noticed in [11], if $\lambda_{i_{0}+1}\leq 0$ as in the Oseledec theorem, then $h^{E}_{\mu}(f)=h_{\mu}(f)$.
When $E$ is $1$-dimensional, we also denote $h^{1}_{\mu}(f)=h^{E}_{\mu}(f)$ as in the statement of Theorem C and Theorem D.
Let $x\in M$, $n\in{\mathbb{N}}$ and $\rho>0$, the usual Bowen ball is defined by
$B(x,n,\rho):=\{y\in M:d(f^{j}(x),f^{j}(y))<\rho,\ \forall 0\leq j<n\}.$
It is clear that for $\mu$-almost every $x$
$V^{E}(x,n,\rho)\subset W^{E}_{\rm loc}(x)\cap B(x,n,\rho),$
and so, we have $\mu_{\xi(x)}(V^{E}(x,n,\rho))\leq\mu_{\xi(x)}(B(x,n,\rho))$.
Proposition 2.1.
For any $\varepsilon>0$ and any $\delta>0$, there exists $K\subset M$ with $\mu(K)>1-\delta$ and $\rho>0$, such that
$\forall x\in K,\quad h_{\mu}^{E}(f)\leq\liminf_{n\rightarrow+\infty}\frac{\log\mu_{\xi(x)}\left(K\cap B(x,n,\rho)\right)}{-n}+\varepsilon.$
Proof.
For any $\varepsilon>0$ and $\delta>0$, choose $\varepsilon^{\prime}\in(0,\varepsilon)$, such that for $n$ large enough, one has that
$\sum_{j=0}^{\lceil n\varepsilon^{\prime}\rceil}\tbinom{n}{j}\leq e^{n\varepsilon/3}.$
Choose a set $F$ with
$\mu(F)>1-\min\{\frac{\varepsilon^{\prime}}{5\|Df\|^{\operatorname{dim}M}_{\sup}},\frac{\delta}{2}\}$
and $\rho_{F}>0$ such that for every $0<\rho<\rho_{F}$ and every $x\in F$
  1. (1)
    $f(V^{E}(f^{-1}(x),1,2\rho))\subset W^{E}_{\operatorname{loc}}(x)$;
  2. (2)
    $d^{E}_{x}\left(B(y,\rho)\cap W^{E}_{\operatorname{loc}}(x)\right)<3\rho$ for any $y\in M$.
By Equation (1), there exists $K\subset F$ with $\mu(K)>1-\delta$, such that
$\exists 0<\rho_{K}<\rho_{F},\ \forall 0<\rho<\rho_{K}/6,~{}\ \exists N:=N(f,\rho)\in{\mathbb{N}}\quad s.t.\quad\forall n\geq N,\ \forall x\in K$
#$\#\{0\leq k<n:f^{k}(x)\notin F\}\leq\min\{\frac{n\varepsilon}{4\|Df\|_{\sup}^{\operatorname{dim}M}},n\varepsilon^{\prime}\},\quad\mu_{\xi(x)}\left(V^{E}(x,n,6\rho)\right)\leq\exp\left(-n(h^{E}_{\mu}(f)+\dfrac{\varepsilon}{3})\right).$
Fix $n\geq N$ large enough, by the choice of $\varepsilon^{\prime}$, one has that
#$\#\left\{E(x):x\in K\right\}\leq\sum_{j=0}^{\lceil n\varepsilon^{\prime}\rceil}\tbinom{n}{j}\leq e^{\frac{n\varepsilon}{3}},$
where $E(x)=\{0\leq k<n:f^{k}(x)\notin F\}$. For a fixed type ${\cal E}\subset[0,n)$, denote by $K_{\cal E}:=\{x\in K,\ E(x)={\cal E}\}$. Then $K$ is covered by at most $e^{\frac{n\varepsilon}{3}}$ disjoint subsets of the form $K_{\cal E}$, and
$\forall x\in K,\quad\mu_{\xi(x)}\left(K\cap B(x,n,\rho)\right)\leq e^{\frac{n\varepsilon}{3}}\sup_{\cal E}\mu_{\xi(x)}(K_{\cal E}\cap B(x,n,\rho)).$
We now fix a type $\cal E$ such that the $\sup$ is obtained.
Claim.
For any $1\leq k\leq n$, there exists $\{x_{1}^{k},\cdots,x_{c(k)}^{k}\}\subset K_{\cal E}$ and sub-manifolds $(D_{k}(x_{j}^{k}))_{j=1}^{c(k)}$ of $W^{E}_{\operatorname{loc}}(x)$, such that
$K_{\cal E}\cap B(x,k,\rho)\cap W^{E}_{\operatorname{loc}}(x)\subset\bigcup_{j=1}^{c(k)}D_{k}(x_{j}^{k});$
$\operatorname{Diam}_{d^{E}_{f^{i}(x)}}\left(f^{i}(D_{k}(x_{j}^{k}))\right)\leq 3\rho,\ \forall 0\leq i<k,\ \forall 1\leq j\leq c(k);$
$c(k)\leq||Df||_{\sup}^{\#\{0\leq i<k:i\in E\}\cdot\operatorname{dim}M}.$
Proof.
For $k=1$, since $K\subset F$ one has $0\notin{\cal E}$. We can choose $x_{1}^{1}\in K_{\cal E}$ such that
$K_{\cal E}\cap B(x,\rho)\cap W^{E}_{\operatorname{loc}}(x_{1})\subset B(x,\rho)\cap W^{E}_{\operatorname{loc}}(x_{1}):=D(x_{1}^{1}),$
and $\operatorname{Diam}_{d^{E}_{x}}(D(x_{1}^{1}))<3\rho$ by item (2).
Assume that the claim is true for $k$, i.e.,
$K_{\cal E}\cap B(x,k,\rho)\cap W^{E}_{\operatorname{loc}}(x)\subset\bigcup_{j=1}^{c(k)}D_{k}(x_{j}^{k}).$
We now prove is also true for $k+1$. Note that
$B(x,k+1,\rho)=f^{-k}(B(f^{k}x,\rho))\bigcap B(x,k,\rho).$
It suffices to show that for each $z=x_{j}^{k}$, there exists $\{y_{k}\}_{k=1}^{m}\subset K_{\cal E}$ and sub-manifolds $(D(f^{k}y_{i}))_{i=1}^{m}$ of $W^{E}_{\operatorname{loc}}(x)$ such that
$f^{k}D_{k}(z)\cap B(f^{k}(x),\rho)\subset f^{k}\left(\bigcup_{j=1}^{m}D_{k+1}(y_{j})\right);$
$\operatorname{Diam}_{d_{f^{i}(x)}^{E}}(f^{i}D_{k+1}(y_{j}))\leq 3\rho,\ \forall 0\leq i\leq k,\ \forall j=1,\cdots,m;$
$\quad m=1,\ k\notin E;\ m\leq||Df||_{\sup}^{\operatorname{dim}M},\ k\in E.$
If $k\notin{\cal E}$, then $f^{k}z\in F$,
$f^{k}(D_{k}(z))\cap B(f^{k}x,\rho)\subset W^{E}_{\operatorname{loc}}(f^{k}z)\cap B(f^{k}x,\rho):=D(f^{k}z),$
and $\operatorname{Diam}_{d^{E}_{f^{k}(x)}}(D(f^{k}z))<3\rho$. Hence, let $y_{1}=z$ and $D_{k+1}(y_{1})=f^{-k}(D(f^{k}z))\cap D_{k}(z)$, we have
$f^{k}D_{k}(z)\cap B(f^{k}(x),\rho)\subset f^{k}D_{k+1}(y_{1}),$
and $\operatorname{Diam}_{d^{E}_{f^{j}(x)}}(f^{j}D_{k+1}(y_{1}))<3\rho$ for every $0\leq i\leq k$.
If $k\in{\cal E}$, then $f^{k}z\notin F$, there exist $\{y_{1},\cdots,y_{m}\}\subset D_{k}(z)$ and sub-manifolds $(D(f^{k}y_{i}))_{i=1}^{m}$ of $f^{k}D_{k}(z)$ with $m\leq||Df||_{\sup}^{\operatorname{dim}M}$ such that
$f^{k}D_{k}(z)\subset\bigcup_{i=1}^{m}D(f^{k}y_{i}),\ \operatorname{Diam}_{d^{E}_{f^{k}(x)}}(D(f^{k}y_{i}))\leq 3\rho,\ \forall i=1,\cdots,m.$
Hence, let $D_{k+1}(y_{i})=f^{-k}(D(f^{k}y_{i}))\cap D_{k}(z)$, we have
$f^{k}D_{k}(z)\cap B(f^{k}(x),\rho)$ $\subset\bigcup_{j=1}^{m}D(f^{k}y_{j})\cap f^{k}(D_{k}(z))$
$\subset f^{k}\left(\bigcup_{j=1}^{m}D_{k+1}(y_{j})\right),$
and $\operatorname{Diam}_{d^{E}_{f^{j}(x)}}(f^{j}D_{k+1}(y_{i}))<3\rho$ for any $0\leq j\leq k$ and any $1\leq i\leq m$. Hence, the claim is proved. ∎
By the claim, we have
$\mu_{\xi(x)}\left(K\cap B(x,n,\rho)\right)$ $\leq e^{\frac{n\varepsilon}{3}}\cdot c(n)\cdot e^{-n(h_{\mu}^{E}(f)-\varepsilon/3)}$
$\leq e^{\frac{n\varepsilon}{3}}\cdot e^{(n\varepsilon/3)\frac{\log||Df||^{\operatorname{dim}M}_{\sup}}{||Df||^{\operatorname{dim}M}_{\sup}}}\cdot e^{-n(h_{\mu}^{E}(f)-\varepsilon/3)}$
$\leq e^{-n(h_{\mu}^{E}(f)-\varepsilon)}.$
This completes the proof. ∎
By Proposition 2.1, for $\varepsilon>0$ we can choose a compact set $K$ with $\mu(K)>1/2$ and $\rho>0$ such that
$\forall x\in K,\quad{\rm h}_{\mu}^{E}(f)\leq\liminf_{n\rightarrow+\infty}-\frac{1}{n}\log\mu_{\xi(x)}\left(K\cap B(x,n,\rho)\right)+\varepsilon.$ (2)
Let $x_{0}\in K,\Sigma\subset W^{E}_{\operatorname{loc}}(x_{0})$ with $\mu_{\xi(x_{0})}(K\cap\Sigma)>0$.
Proposition 2.2.
Let $\varepsilon$, $\rho$, $K$ and $\Sigma$ be chosen above, for any finite partition $\cal P$ satisfying ${\rm Diam}({\cal P})<\rho$ one has
${\rm h}^{E}_{\mu}(f)\leq\liminf_{n\to\infty}\frac{1}{n}\log\#\{P\cap K\cap\Sigma\neq\emptyset,~{}P\in{\cal P}^{n}\}+\varepsilon.$
Proof.
We consider the probability measure $\mu_{\xi(x_{0}),K}$ by the following way:
$\forall~{}\textrm{Borel set}~{}A,~{}~{}~{}\mu_{\xi(x_{0}),K}(A):=\mu_{\xi(x_{0}),K}^{\Sigma}(A)=\frac{\mu_{\xi(x_{0})}(\Sigma\cap K\cap A)}{\mu_{\xi(x_{0})}(K\cap\Sigma)}.$
By the definition of the static entropy $H$, one has
$H_{\mu_{\xi(x_{0}),K}}({\cal P}^{n})\leq\log\#\{P\cap K\cap\Sigma\neq\emptyset,~{}P\in{\cal P}^{n}\}.$
It suffices to show that
${\rm h}_{\mu}(f)\leq\liminf_{n\to\infty}\frac{1}{n}H_{\mu_{\xi(x_{0}),K}}({\cal P}^{n})+\varepsilon.$
Note that
$\liminf_{n\rightarrow+\infty}\frac{1}{n}H_{\mu_{\xi(x_{0}),K}}({\cal P}^{n})$ $=\liminf_{n\rightarrow+\infty}\int-\frac{1}{n}\log\mu_{\xi(x_{0}),K}({\cal P}^{n}(y))\ {\rm d}\mu_{\xi(x_{0}),K}(y),\ \ \text{by the definition}$
[…]
By the definition of $\mu_{\xi(x_{0}),K}$, one has
$\mu_{\xi(x_{0}),K}({\cal P}^{n}(y))=\dfrac{\mu_{\xi(x_{0})}({\cal P}^{n}(y)\cap K\cap\Sigma)}{\mu_{\xi(x_{0})}(K\cap\Sigma)}.$
For any $y\in K\cap\xi(x_{0})$, we have
$\mu_{\xi(x_{0})}({\cal P}^{n}(y)\cap K\cap\Sigma)\leq\mu_{\xi(y)}({\cal P}^{n}(y)\cap K)\leq\mu_{\xi(y)}(B(y,n,\rho)\cap K).$
By Equation (2) and the measure $\mu_{\xi(x_{0}),K}$ supported on $K\cap\xi(x_{0})$ one has
$\liminf_{n\rightarrow\infty}-\frac{1}{n}\log\mu_{\xi(x_{0}),K}({\cal P}^{n}(y))\geq{\rm h}_{\mu}(f)-\varepsilon\quad\text{for}\ \mu_{\xi(x_{0}),K}\text{-}a.e.\ y.$
Therefore, we have
$\liminf_{n\rightarrow\infty}\frac{1}{n}H_{\mu_{\xi(x_{0}),K}}({\cal P}^{n})\geq{\rm h}_{\mu}(f)-\varepsilon.$
This completes the proof. ∎
3 Yomdin-Burguet’s reparametrization lemma
We need the notion of bounded curves by Burguet [2, 3].
A curve in $M$ is identical to a $C^{1}$ map $\sigma:~{}[-1,1]\to M$. Denote by $\sigma_{*}$ the image of $\sigma$, i.e., $\sigma_{*}=\sigma([-1,1])\subset M$. Given $L>0$, a $C^{r}$-curve $\sigma$ is said to be a $1/L$-curve, if
$\max_{s=2,\cdots,r}\|D^{s}\sigma\|_{\rm sup}\leq\frac{1}{L}\|D\sigma\|_{\rm sup}.$
Given $L>0$ and $\varepsilon>0$, a $C^{r}$ $1/L$-curve is said to be $\varepsilon$-bounded, if $\|D\sigma\|_{\rm sup}\leq\varepsilon$.
For a linear normed space $A$, for $\delta>0$, we use the notation
$A(\delta):=\{v\in A:~{}\|v\|<\delta\}.$
Since $M$ is a compact manifold, one can choose $r(M)>0$ and $\rho(M)>0$ such that for any $x\in M$, the exponential map $\exp_{x}$ is injective on $T_{x}M(r(M))$, and $\exp_{x}(T_{x}M(r(M)))\supset B(x,\rho(M))$.
A curve $\sigma$ is said to be essentially $1/3$-graph, if there is $x\in\sigma_{*}$, there is a linear subspace $E\subset T_{x}M$ and a subset $E_{\sigma}\subset E$ containing 00, there is a $C^{1}$ map $\varphi:~{}E\to E^{\perp}$ whose Lipschitz constant is less than $1/3$, such that $\varphi(0)=0$, $\sigma_{*}\subset B(x,\rho(M))$ and $\exp_{x}^{-1}\sigma_{*}=\{(v,\varphi(v)):~{}v\in E_{\sigma}\}$. An essential $1/3$-graph is $\beta$-large, if in the above definition, one has that $E_{\sigma}\supset E(\beta)$.
Lemma 3.1.
There is $L>0$ such that for any $1/L$-curve $\sigma$, if ${\rm diam}(\sigma_{*})<\rho(M)$ and $\|D\sigma(0)\|<1/L$, then $\sigma$ is an essentially $1/3$-graph.
Proof.
Take $x=\sigma(0)$. Since ${\rm diam}(\sigma_{*})<\rho(M)$, one can consider the curve $\exp_{x}^{-1}\circ\sigma$ in $T_{x}M$. By taking $E=T_{x}\sigma_{*}$ one can represent it as a graph of a map $\psi$ from a subset of $E$ to $E^{\perp}$.
By following the proof of [3, Section 4.1], one know that the tangent space of every point $\exp_{x}^{-1}\circ\sigma(t)$ is close to $E$. Thus we know the Lipchitz constant of $\psi$ is less than $1/3$. ∎
Fix $L$ as in Lemma 3.1. Then we simply call $1/L$-curves by bounded curves by following Burguet [3]. Note that in Burguet [3], $L$ is taken to be $6$. We have to adapt the notion a bit since we want to control the Lipschitz constant precisely.
Lemma 3.2.
Given $\varepsilon>0$, there is $\beta_{\varepsilon}>0$ such that if $\sigma$ is a bounded curve and $\|D\sigma(t)\|\geq\varepsilon$ for some $t\in[-1,1]$, then $\sigma$ is $2\beta_{\varepsilon}$-large.
Proof.
By following the proof of [3, Section 5.4], the lengths of $\sigma([-1,0])$ and $\sigma([0,1])$ are both larger than $\delta$ for some constant $\delta$ related to $\varepsilon$. After representing it as the graph of a map $\psi$ in the tangent space, by Lemma 3.1, the Lipschitz constant of $\psi$ is less than $1/3$. Thus, one knows that the domain of $\psi$ containing $E(\beta)$ for some $\beta>0$ related to $\delta$, hence related to $\varepsilon$. ∎
A map $\theta:~{}[-1,1]\to[-1,1]$ is also called a reparametrization. A reparametrization $\theta$ is said to be affine if $\theta^{\prime}$ is constant and positive. Note that an affine reparametrization must be contracting, i.e., $\theta^{\prime}\leq 1$.
One has the following reparametrization lemma from Burguet [4, Lemma12]. The statement is a bit different, but the proof is essentially the same.
Lemma 3.3.
Given $r\geq 2$, there is $C_{r}>0$ with the following property.
For any $C^{r}$ diffeomorphism $g:~{}M\to M$, there exists $\varepsilon_{g}^{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{g}^{0})$, for any $\varepsilon$-strongly bounded $C^{r}$ curve, for any two integers $k$ and $k^{\prime}$, there is a family $\Theta$ of affine reparametrizations, such that
  1. 1.
    $\{t\in[-1,1]:~{}x=\sigma(t),\lceil\log\|Dg(x)\|\rceil=k,~{}\lceil\log\|Dg|_{T_{x}{\sigma_{*}}}\|\rceil=k^{\prime}\}\subset\bigcup_{\theta\in\Theta}\theta([-1,1])$
  2. 2.
    $g\circ\sigma\circ\theta$ is bounded;
  3. 3.
    #$\#\Theta\leq C_{r}{\rm e}^{\frac{k-k^{\prime}}{r-1}}$.
Lemma 3.4.
For any $\varepsilon>0$, for any diffeomorphism $g$, if $\sigma$ is $\varepsilon$-bounded, $g\circ\sigma$ is bounded, then there is a family $\Theta_{g}$ of affine contracting reparametrizations, such that
#$\#\Theta_{g}\leq\|Dg\|_{\rm sup}+2,$
and for any $\theta\in\Theta_{g}$, $g\circ\sigma\circ\theta$ is $\varepsilon$-bounded.
Proof.
One can take a family $\Theta_{g}$ of affine reparametrizations such that #$\#\Theta_{g}\leq\lceil\|Dg\|_{\rm sup}\rceil\leq\|Dg\|_{\rm sup}+2$ such that for any $\theta\in\Theta_{g}$, one has $\theta^{\prime}\leq 1/\|Dg\|_{\rm sup}$ and
$\bigcup_{\theta\in\Theta_{g}}\theta([-1,1])=[-1,1].$
We have to check that $g\circ\sigma\circ\theta$ is $\varepsilon$-bounded. Indeed,
$\|D(g\circ\sigma\circ\theta)\|\leq\|Dg\|_{\rm sup}\cdot\|D\sigma\|\cdot\|D\theta\|\leq\|Dg\|_{\rm sup}\cdot\varepsilon\cdot\frac{1}{\|Dg\|_{\rm sup}}\leq\varepsilon.$
4 Choose constants and sets
Let $\alpha>0$ be as in the statement of Theorem C.
Choose $q$.
$q$ is a large integer such that
  • $\frac{1}{q}\log\|Df^{q}\|_{\rm sup}-R(f)<\alpha/10.$ (3)
  • $\frac{1}{q}\log q<\alpha/10.$ (4)
  • $q\alpha>\frac{1}{2r}.$ (5)
  • $\frac{1}{q}\log\big{(}\max\{\log\|Df\|_{\rm sup},~{}\log\|Df^{-1}\|_{\rm sup}\}+2\big{)}<\frac{\alpha}{10}.$ (6)
  • For the constant $C_{r}$ in Lemma 3.3.
    $\frac{1}{q}\log(2C_{r})<\alpha/10.$ (7)
Choose $\beta$
After $q$ is chosen, for the map $g=f^{q}$, one has the constant $\varepsilon^{0}_{f^{q}}$ from Lemma 3.3. By taking $\varepsilon_{g}\in(0,\varepsilon^{0}_{f^{q}})$, one assumes that
$\|Df\|_{\rm sup}^{q}\varepsilon_{g}\leq 1$ (8)
Choose $\beta>0$ such that
  • $\beta<\beta_{\varepsilon_{g}}$ as in Lemma 3.2.
  • $\max_{1\leq j\leq q}\{\|Df^{j}\|_{\rm sup}\}\cdot\beta\leq\varepsilon_{f^{q}}.$
Choose $c$.
As in Katok’s paper [9, Equation 1.3], by using Stirling’s formula
$\forall r<\frac{1}{2},\ \forall k\geq 1,\quad\lim_{n\to\infty}\frac{1}{n}\log\sum_{j=0}^{\lceil nr\rceil}\tbinom{n}{j}\cdot k^{j}=r\log k-r\log r-(1-r)\log(1-r).$
We first choose $\alpha_{1}>0$ such that there exists $N(\alpha_{1})\in{\mathbb{N}}$ for which
$\forall n\geq N(\alpha_{1}),~{}\sum_{j=0}^{\lceil n\alpha_{1}\rceil}\tbinom{n}{j}\leq e^{n\alpha/10}~{}\text{and}~{}\frac{\alpha_{1}}{q}\log\|Df^{q}\|_{\rm sup}\leq\alpha/10.$ (9)
We then choose $c>0$ by taking
$c=\frac{\alpha_{1}}{4q}.$ (10)
Assume that $\mu$ is an ergodic measure with one dominated Lyapunov exponent. We choose a compact set $K:=K_{\mu}$ satisfying $\mu(K)>1/2$ such that the following properties hold:
  1. 1.
    The limit measures for points in $K$ converges to $\mu$ uniformly: for any $\varepsilon>0$, there is $N\in{\mathbb{N}}$ such that for any $n>N$, for any $x\in K$, one has that $d(\frac{1}{n}\sum_{j=0}^{n-1}\delta_{f^{i}(x)},\mu)<\varepsilon$.
  2. 2.
    For $n$ large enough, one has that
    $\forall x\in K,\quad\#\{0\leq i<n,f^{i}(x)\in L^{u,1}(\beta)\}<n(\mu(L^{u,1}(\beta))+c).$ (11)
  3. 3.
    For $m$ large enough,
    $\frac{1}{m}\sum_{j=0}^{m-1}\log\|Df^{q}|_{E^{u,1}(f^{jq}(y))}\|\geq\lambda_{1}(\mu,f^{q})-2q\alpha/5$
  4. 4.
    For $\alpha>0$, there is $\rho>0$ such that for any $x\in K$, one has that
    $h_{\mu}^{1}(f)\leq\liminf_{n\to\infty}-\frac{1}{n}\log(\mu_{\xi(x)}(B(x,n,\rho)\cap K))+\alpha/10.$
We can choose a point $x_{0}\in K$ such that
  • $x_{0}\in K$, $\mu_{\xi(x_{0})}(K)>0$ and for any $\delta>0$ small enough, one has that $\mu_{\xi(x_{0})}(K\cap W^{u}_{\delta}(x_{0}))>0$.
5 Cover $1$-dimensional unstable curves
Assume that $\mu$ is an ergodic measure with one dominated Lyapunov exponent. Recall the constants chosen as in Section 4, the compact set $K$ and $x_{0}\in K$ chosen as in Section 4. We take a reparametrization $\sigma:~{}[-1,1]\to W^{1}_{\rm loc}(x_{0})$ such that $\sigma(0)=x_{0}$ and $\sigma$ is $\varepsilon_{f^{q}}$-bounded.
Theorem 5.1.
Assume that $\mu(L^{u,1}(\beta))\leq c$. For any $n\in{\mathbb{N}}$, there is a family of affine reparametrizations $\Gamma_{n}$ having the following properties:
  1. 1.
    $\limsup_{n\to\infty}\frac{1}{n}\log\#\Gamma_{n}\leq\frac{1}{r}R(f)+\alpha/2$
  2. 2.
    $\bigcup_{\gamma\in\Gamma_{n}}\sigma\circ\gamma([-1,1])\supset K\cap\sigma_{*}$;
  3. 3.
    $\|D(f^{j}\circ\sigma\circ\gamma)\|\leq 1,~{}~{}\forall j=0,1,\cdots,n-1$.
For the reparametrizations $\{\Gamma_{n}\}$ in Theorem 5.1, for a partition $\mathcal{P}$ whose diameter is less than $\rho$ as in Proposition 2.2, we define
$D(n):=\sup_{\gamma\in\Gamma_{n}}\#\{P:~{}P\cap K\cap(\sigma\circ\gamma)_{*}\neq\emptyset,~{}P\in{\cal P}^{n}\}.$
Inspired by [4, Page 1498], one has the following proposition.
Proposition 5.2.
Assume that $\mu(\partial{\cal P})=0$,
$\lim_{n\to\infty}\frac{1}{n}\log D(n)=0.$
Proof.
We will prove by contradiction. Assume the conclusion is not true. Then there is $\kappa>0$, such that
$\limsup_{n\to\infty}\frac{1}{n}\log D(n)>4\kappa\log\#{\cal P}.$ (12)
Since $\mu(\partial{\mathcal{P}})=0$, there is $\chi\in(0,1)$ small enough, such that
$\mu(B(\partial{\cal P},\chi))<\kappa.$ (13)
Equation (12) implies that there are sufficiently large $n$, and $\gamma\in\Gamma_{n}$ such that
#$\#\{P:~{}P\cap K\cap(\sigma\circ\gamma)_{*}\neq\emptyset,~{}P\in{\cal P}^{n}\}>4([1/\chi]+2)(\#{\cal P})^{2n\kappa}.$ (14)
Equation (13) implies that for any point $x\in K$, one has that for $n$ large enough,
#$\#\{0\leq j\leq n-1:~{}f^{j}(x)\in B(\partial{\cal P},\chi)\}<n\kappa.$
Now we take a family of reparametrizations $\Theta$ such that
  • #$\#\Theta<4([1/\chi]+2)$;
  • $\big{|}\theta^{\prime}\big{|}<\chi$, for any $\theta\in\Theta$;
  • $\bigcup_{\theta\in\Theta}\theta([-1,1])=[-1,1]$.
Thus, there is $\theta\in\Theta$, such that
#$\#\{P:~{}P\cap K\cap(\sigma\circ\gamma\circ\theta)_{*}\neq\emptyset,~{}P\in{\cal P}^{n}\}\geq(\#{\cal P})^{2n\kappa}.$
On the other hand, for any $\theta\in\Theta$, one always has
#$~{}~{}~{}~{}~{}~{}\#\{P:~{}P\cap K\cap(\sigma\circ\gamma\circ\theta)_{*}\neq\emptyset,~{}P\in{\cal P}^{n}\}$
$\leq\prod_{j=0}^{n-1}\#\{P\in{\cal P}:~{}P\cap f^{j}(\sigma\circ\gamma\circ\theta([-1,1]))\cap f^{j}(K)\neq\emptyset\}$
Since $\|D(f^{j}\circ\sigma\circ\gamma)\|\leq 1,~{}~{}\forall j=0,1,\cdots,n-1$, one has that $\|D(f^{j}\circ\sigma\circ\gamma\circ\theta)\|\leq\chi/2,~{}~{}\forall j=0,1,\cdots,n-1$. This implies that the diameter of $f^{j}\circ\sigma\circ\gamma\circ\theta([-1,1])$ is less than $\chi$. Thus,
  • #$\#\{P\in{\cal P}:~{}P\cap f^{j}(\sigma\circ\gamma\circ\theta([-1,1]))\cap f^{j}(K)\neq\emptyset\}>1$ if and only if $f^{j}\circ\sigma\circ\gamma\circ\theta([-1,1])\cap f^{j}(K)$ has non-empty intersection with $B(\partial{\cal P},\chi)$.
Thus, we have
$\prod_{j=0}^{n-1}\#\{P\in{\cal P}:~{}P\cap f^{j}(\sigma\circ\gamma\circ\theta([-1,1]))\cap f^{j}(K)\neq\emptyset\}\leq\big{(}\#{\cal P}\big{)}^{\sup_{x\in K}\#\{j:~{}f^{j}(x)\in B(\partial{\cal P},\chi)\}}\leq\big{(}\#{\cal P}\big{)}^{n\kappa}.$
This contradicts to Equation (14).
6 Proof of Theorem C
For $\alpha>0$ as in the statement of Theorem C. We have chosen $\beta>0$ and $c>0$ as in Section 4. Let $\mu$ be an ergodic measure as in the statement of Theorem C, i.e.,
$h^{1}_{\mu}(f)>\alpha+\frac{R(f)}{r}.$
We will prove Theorem C by absurd, i.e., we assume that $\mu(L^{u}(\beta))\leq c$.
We choose the compact set $K$ and the constant $\rho$ as in Section 4. Choose $x_{0}\in K$ and a reparametrization $\sigma:~{}[-1,1]\to W^{1}_{\rm loc}(x_{0})$ is $\varepsilon_{g}$-bounded.
We choose a partition $\mathcal{P}$ such that ${\rm Diam}(\mathcal{P})<\rho$ and $\mu(\partial{\mathcal{P}})=0$.
Recall the family of reparametrizations $\Gamma_{n}$ in Theorem 5.1 and $D(n)$ in Proposition 5.2.
Since we have,
#$\#\{P:~{}P\cap K\cap\sigma_{*}\neq\emptyset,~{}P\in{\cal P}^{n}\}\leq\#\Gamma_{n}\sup_{\gamma\in\Gamma_{n}}\{P:~{}P\cap K\cap(\sigma\circ\gamma)_{*}\neq\emptyset,~{}P\in{\cal P}^{n}\}$
By Proposition 2.2, we have that
$h_{\mu}^{1}(f)\leq\liminf_{n\to\infty}\frac{1}{n}\log\#\Gamma_{n}+\lim_{n\to\infty}\frac{1}{n}\log D(n)<\frac{1}{r}R(f)+\alpha$
by using Theorem 5.1 and Proposition 5.2. Now we have a contradiction to the assumption of Theorem C.
7 The proof of Theorem 5.1
For $n$ in Theorem 5.1, we write $n=mq+\ell$, where $\ell\in[0,q)$. In this section, we always take $g=f^{q}$.
Recall the relationship of $c$ and $\alpha_{1}$ from Equation (10).
Lemma 7.1.
For any $x\in K$, for $m$ large enough, one has that
#$\#\{0\leq i\leq m-1:~{}f^{qi}(x)\in L^{u,1}(\beta)\}\leq\alpha_{1}m.$
Proof.
Since $\mu(L^{u,1}(\beta))<c$, one has that for $m$ large enough, for any $x\in K$, by Equation (11)
#$\#\{0\leq j\leq mq-1:~{}f^{j}(x)\in L^{u,1}(\beta)\}\leq cmq.$
Clearly, #$\#\{0\leq i\leq m-1:~{}f^{qi}(x)\in L^{u,1}(\beta)\}\leq\#\{0\leq j\leq mq-1:~{}f^{j}(x)\in L^{u,1}(\beta)\}$. Thus, one can conclude by Equation (10). ∎
We define the type
${\cal S}_{m}=\{E\subset[0,m):~{}d_{m}(E)<\alpha_{1}\}.$
By the choice of $\alpha_{1}$ (Equation (9)), for $m$ large enough one has that
$\frac{1}{m}\log(\sum_{j=0}^{\lceil m\alpha_{1}\rceil}\tbinom{m}{j})<\alpha/10,$
One has that
#$\#{\cal S}_{m}\leq{\rm e}^{\alpha n/10}.$ (15)
For any $E\in{\cal S}_{m}$, we consider
$K_{E}=\{x\in K\cap\sigma_{*}:~{}f^{jq}(x)\in L^{u,1}(\beta)~{}{\textrm{iff}}~{}j\in E\}.$
Given $2m$-integers $(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})$, we define a subset of $K_{E}$:
$\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})=\Sigma(E;k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})$
by the following way: $x\in\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})$ if and only if $x\in K_{E}$ and
$\forall 0\leq j\leq m-1,~{}~{}~{}\lceil\log\|Dg(g^{j}(x))\|\rceil=k_{j},~{}\lceil\log\|Dg|_{T_{g^{j}(x)}{(g^{j}(\sigma))_{*}}}\|\rceil=k_{j}^{\prime}$
If $\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})\neq\emptyset$, then the $2m$-tuple $(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})$ is said to be admissible.
Meanwhile, for a point $y\in\sigma_{*}$, for any $j\geq 0$, we define functions:
$k_{j}(y):=\lceil\log\|Dg(g^{j}(y))\|\rceil,~{}~{}~{}k_{j}^{\prime}(y):=\lceil\log\|Dg|_{T_{g^{j}(y)}{(g^{j}(\sigma))_{*}}}\|\rceil.$
Lemma 7.2.
The are at most
$\big{(}q\cdot\max\{\log\|Df\|_{\rm sup}+2,~{}\log\|Df^{-1}\|_{\rm sup}+2\}\big{)}^{2m}$
admissible $2m$-tuples.
Proof.
For any $j$, $k_{j}$ or $k_{j}^{\prime}$ has $\max\{\lceil\log\|Df^{q}\|_{\rm sup}\rceil,\lceil\log\|Df^{-q}\|_{\rm sup}\rceil\}+1$ possibilities. It is less than $q\cdot\max\{\log\|Df\|_{\rm sup}+2,~{}\log\|Df^{-1}\|_{\rm sup}+2\}$. Since we have $2m$ integers, one gets the estimates. ∎
Lemma 7.3.
For $n=mq$ large enough, for $y\in K$, one has that
$\frac{1}{n}\frac{1}{r-1}\sum_{j=0}^{m-1}(k_{j}(y)-k_{j}^{\prime}(y))\leq\frac{1}{r}\frac{1}{q}\log||Df^{q}||_{\rm sup}\leq\frac{1}{r}R(f)+\alpha/10$
Proof.
By our assumption:
$h^{1}_{\mu}(f)>\alpha+\frac{R(f)}{r},$
Thus, by Equation (3),
$h^{1}_{\mu}(f^{q})=qh^{1}_{\mu}(f)\geq q(\frac{R(f)}{r}+\alpha)\geq q\left(\frac{1}{r}(\frac{\log\|Df^{q}\|_{\rm sup}}{q}-\frac{\alpha}{10})+\alpha\right)\geq\frac{\log\|Df^{q}\|_{\rm sup}}{r}+9q\alpha/10.$
By the Ruelle inequality [18] (see [11, Section 10.1] for the version of partial entropy), one has that $\lambda_{1}(\mu,f^{q})\geq h^{1}_{\mu}(f^{q})>\frac{1}{r}\log\|Df^{q}\|_{\rm sup}+9q\alpha/10$. Thus, for $m$ large enough, for any $y\in K$,
$\frac{1}{m}\sum_{j=0}^{m-1}\log\|Df^{q}|_{E^{u,1}(f^{jq}(y))}\|-\frac{1}{r}\frac{1}{m}\sum_{j=0}^{m-1}\log\|Df^{q}(f^{jq}(y))\|$
$\geq\lambda_{1}(\mu,f^{q})-2q\alpha/5-\frac{1}{r}\log\|Df^{q}\|_{\rm sup}>q\cdot\alpha/2.$
Together with Equation (5), this implies that
$\sum_{j=0}^{m-1}k_{j}^{\prime}(y)-\frac{1}{r}\sum_{j=0}^{m-1}k_{j}(y)>m(q\cdot\alpha/2-1/r)>0,$
in other words,
$\sum_{j=0}^{m-1}k_{j}^{\prime}(y)>\frac{1}{r}\sum_{j=0}^{m-1}k_{j}(y).$
Thus, one has
$\frac{1}{r-1}\sum_{j=0}^{m-1}(k_{j}(y)-k_{j}^{\prime}(y))\leq\frac{1}{r-1}\frac{r-1}{r}\sum_{j=0}^{m-1}k_{j}(y)\leq\frac{1}{r}m\log\|Df^{q}\|_{\rm sup}.$
By dividing $n=mq$, by Equation (3), one has that
$\frac{1}{n}\frac{1}{r-1}\sum_{j=0}^{m-1}(k_{j}(y)-k_{j}^{\prime}(y))\leq\frac{1}{r}\frac{1}{q}\log\|Df^{q}\|_{\rm sup}\leq\frac{1}{r}R(f)+\alpha/10.$
We will prove the following proposition.
Proposition 7.4.
For the set $\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime})$, for any $1\leq j\leq m$, there is a family $\Gamma_{j}^{\prime}=\Gamma_{j}^{\prime}(\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime}))$ of reparametrizations satisfying
#$\#\Gamma_{j}^{\prime}\leq 2^{j}C_{r}^{j}\exp\{\sum_{i=0}^{j-1}\frac{k_{i}-k_{i}^{\prime}}{r-1}\}(\|Dg\|_{\rm sup}+2)^{\#\{0<i\leq j,~{}i\in E\}},$
and one has the following properties:
  1. 1.
    $\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime})\subset\bigcup_{\gamma\in\Gamma_{j}^{\prime}}\sigma\circ\gamma([-1,1])$
  2. 2.
    for any $\gamma\in\Gamma_{j}^{\prime}$, $g^{i}\circ\sigma\circ\gamma$ is $\varepsilon_{g}$-bounded for any $1\leq i\leq j$.
Proof.
We will prove this proposition by induction. We first consider the case $j=1$. In this case, we consider the set $\Sigma(k_{0},k_{0}^{\prime})$. By applying the reparametrization Lemma (Lemma 3.3), a family $\Theta$ of affine reparametrizations, whose cardinality satisfying
#$\#\Theta\leq C_{r}{\rm e}^{\frac{k_{0}-k_{0}^{\prime}}{r-1}},$
such that
  1. 1.
    $\Sigma(k_{0},k_{0}^{\prime})\subset\bigcup_{\theta\in\Theta}\sigma\circ\theta([-1,1])$
  2. 2.
    $g\circ\sigma\circ\theta$ is bounded for any $\theta\in\Theta$;
Note that $g\circ\sigma\circ\theta$ may not be $\varepsilon_{g}$-bounded. However, one notices that $\sigma\circ\theta$ is $\varepsilon_{g}$-bounded.
We have two cases
  • $1\in E$: this means $g(x)=f^{q}(x)\in L^{u,1}(\beta)$ for all $x\in\Sigma(k_{0},k_{0}^{\prime})$.
  • $1\notin E$: this means $g(x)=f^{q}(x)\notin L^{u,1}(\beta)$ for all $x\in\Sigma(k_{0},k_{0}^{\prime})$.
For the case $1\in E$, one just applies Lemma 3.4 to have a new family $\Theta_{g}$ of affine reparametrizations such that $g\circ\sigma\circ\theta\circ\theta_{g}$ is $\varepsilon_{g}$-bounded. Define
$\Gamma_{1}^{\prime}:=\{\gamma=\theta\circ\theta_{g}:~{}\theta\in\Theta,~{}\theta_{g}\in\Theta_{g}\}.$
It is clear that
#$\#\Gamma_{1}^{\prime}\leq\#\Theta\cdot\#\Theta_{g}\leq C_{r}{\rm e}^{\frac{k-k^{\prime}}{r-1}}(\|Dg\|_{\rm sup}+2).$
For the case $1\notin E$, for any $\theta\in\Theta$, there are two cases. Either $g\circ\sigma\circ\theta$ is $\varepsilon_{g}$-bounded, or $g\circ\sigma\circ\theta$ is not $\varepsilon_{g}$-bounded.
If $g\circ\sigma\circ\theta$ is not $\varepsilon_{g}$-bounded, then by Lemma 3.2, the length of $g\circ\sigma\circ\theta$ is larger than $2\beta_{\varepsilon_{g}}$ One defines two affine reparametrizations $\gamma_{\theta}^{-1}$ and $\gamma_{\theta}^{1}$ such that
  • $\gamma_{\theta}^{-1}(-1)=-1$ and the length of $g\circ\sigma\circ\theta\circ\gamma_{\theta}^{-1}$ is $\beta_{\varepsilon_{g}}$.
  • $\gamma_{\theta}^{1}(1)=1$ and the length of $g\circ\sigma\circ\theta\circ\gamma_{\theta}^{1}$ is $\beta_{\varepsilon_{g}}$.
Claim.
For any $x\in\Sigma(k_{0},k_{0}^{\prime})$ and $g(x)\in g\circ\sigma\circ\theta([-1,1])$, one has
$x\in\sigma\circ\theta\circ\gamma_{\theta}^{-1}([-1,1])\cup\sigma\circ\theta\circ\gamma_{\theta}^{1}([-1,1]).$
Proof of the claim.
Otherwise, one will have $g(x)\notin g\circ\sigma\circ\theta\circ\gamma_{\theta}^{-1}([-1,1])\cup g\circ\sigma\circ\theta\circ\gamma_{\theta}^{1}([-1,1])$. But by Lemma 3.2, the lengths of $g\circ\sigma\circ\theta\circ\gamma_{\theta}^{-1}([-1,1])$ and $g\circ\sigma\circ\theta\circ\gamma_{\theta}^{1}([-1,1])$ are both much larger than $\beta$ since $\beta$ can be chosen to be much smaller than $\beta_{\varepsilon_{g}}$. Note also the image of $g\circ\sigma\circ\theta$ is contained in the unstable manifold of $W^{u}(g(x))$. This implies that $g(x)\in L^{u,1}(\beta)$. Thus, we get a contradiction. ∎
Thus, we have the decomposition $\Theta=\Theta_{L}\cup\Theta_{S}$, where $\theta\in\Theta_{L}$ if and only if $g\circ\sigma\circ\theta$ is not $\varepsilon_{g}$-bounded. Define
$\Theta_{L}^{\pm 1}=\{\theta\circ\gamma_{\theta}^{\pm 1}:~{}\theta\in\Theta_{L}\},$
and
$\Gamma_{1}^{\prime}=\Theta_{L}^{1}\cup\Theta_{L}^{-1}\cup\Theta_{S}.$
Thus for any $\gamma\in\Gamma_{1}^{\prime}$, $g\circ\sigma\circ\gamma$ is $\varepsilon_{g}$-bounded. Indeed, we only have to consider the case of $\Theta_{L}^{\pm 1}$. Since the length is less than $2\beta_{\varepsilon_{g}}$, one knows the $\varepsilon_{g}$-bounded property by Lemma 3.2. From the Claim above, it is clear that
$\Sigma(k_{0},k_{0}^{\prime})\subset\bigcup_{\gamma\in\Gamma_{1}^{\prime}}\sigma\circ\gamma([-1,1]).$
The cardinality of $\Gamma_{1}^{\prime}$ can be estimated above by the case $1\in E$ or not. Thus, one gets the case of $j=1$.
Assume the case of $j$ is proved. This means one has a family $\Gamma_{j}^{\prime}$ of reparametrizations such that the conclusion holds for $j$. Now we prove the case of $j+1$. In fact it is close to the case $j=1$. For completeness, we give the proof.
Now we consider the set $\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime},k_{j},k_{j}^{\prime})$. Take $\gamma^{\prime}\in\Gamma_{j}^{\prime}$. We know that $g^{i}\circ\sigma\circ\gamma^{\prime}$ is $\varepsilon_{g}$-bounded for any $1\leq i\leq j$.
By applying the reparametrization Lemma (Lemma 3.3), a family $\Theta_{\gamma^{\prime}}$ of affine reparametrizations, whose cardinality satisfying
#$\#\Theta_{\gamma^{\prime}}\leq C_{r}{\rm e}^{\frac{k_{j}-k_{j}^{\prime}}{r-1}},$
such that
  1. 1.
    $\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime},k_{j},k_{j}^{\prime})\cap\sigma\circ\gamma^{\prime}([-1,1])\subset\bigcup_{\theta\in\Theta_{\gamma^{\prime}}}\sigma\circ\gamma^{\prime}\circ\theta([-1,1])$
  2. 2.
    $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is bounded;
Note that $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ may not be $\varepsilon_{g}$-bounded. However, one notices that $g^{j}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is $\varepsilon_{g}$-bounded.
We have two cases
  • $j+1\in E$: this means for all $x\in\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime},k_{j},k_{j}^{\prime})$,$g^{j+1}(x)=f^{(j+1)q}(x)\in L^{u,1}(\beta)$.
  • $j+1\notin E$: this means for all $x\in\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime},k_{j},k_{j}^{\prime})$$g^{j+1}(x)=f^{(j+1)q}(x)\notin L^{u,1}(\beta)$.
For the case $j+1\in E$, one just applies Lemma 3.4 to have a new family $\Theta_{g}$ of affine reparametrizations such that $g^{j}\circ\sigma\circ\gamma^{\prime}\circ\theta\circ\theta_{g}$ is $\varepsilon_{g}$-bounded. Define
$\Gamma_{j+1}^{\prime}:=\{\gamma=\gamma^{\prime}\circ\theta\circ\theta_{g}:~{}\gamma^{\prime}\in\Gamma_{j}^{\prime},~{}\theta\in\Theta,~{}\theta_{g}\in\Theta_{g}\}.$
It is clear that
#$\#\Gamma_{j+1}^{\prime}$$\leq\#\Gamma_{j}^{\prime}\cdot\#\Theta\cdot\#\Theta_{g}\leq 2^{j}C_{r}^{j}\exp\{\sum_{i=0}^{j-1}\frac{k_{i}-k_{i}^{\prime}}{r-1}\}(\|Dg\|_{\rm sup}+2)^{\#\{0<i\leq j,~{}i\in E\}}\cdot C_{r}{\rm e}^{\frac{k_{j}-k_{j}^{\prime}}{r-1}}(\|Dg\|_{\rm sup}+2)$
$\leq 2^{j+1}C_{r}^{j+1}\exp\{\sum_{i=0}^{j}\frac{k_{i}-k_{i}^{\prime}}{r-1}\}(\|Dg\|_{\rm sup}+2)^{\#\{0<i\leq j+1,~{}i\in E\}}$
For the case $j+1\notin E$, for any $\theta\in\Theta$, there are two cases. Either $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is $\varepsilon_{g}$-bounded, or $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is not $\varepsilon_{g}$-bounded.
If $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is not $\varepsilon_{g}$-bounded, one defines two affine reparametrizations $\gamma_{\theta}^{-1}$ and $\gamma_{\theta}^{1}$ such that
  • $\gamma_{\theta}^{-1}(-1)=-1$ and the length $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta\circ\gamma_{\theta}^{-1}$ is $\beta_{\varepsilon_{g}}$.
  • $\gamma_{\theta}^{1}(1)=1$ and the length $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta\circ\gamma_{\theta}^{-1}$ is $\beta_{\varepsilon_{g}}$.
Claim.
For any $x\in\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime},k_{j},k_{j}^{\prime})$ and $g^{j+1}(x)\in g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta([-1,1])$, one has
$x\in\sigma\circ\gamma^{\prime}\circ\theta\circ\gamma_{\theta}^{-1}([-1,1])\cup\sigma\circ\gamma^{\prime}\circ\theta\circ\gamma_{\theta}^{1}([-1,1]).$
Proof of the claim.
The proof is similar to the claim as the case $j=1$. Otherwise, one will have $g^{j+1}(x)\notin g^{j+1}\circ\sigma\circ\theta\circ\gamma_{\theta}^{-1}([-1,1])\cup g^{j+1}\circ\sigma\circ\theta\circ\gamma_{\theta}^{1}([-1,1])$. But by Lemma 3.2, the lengths of $g^{j+1}\circ\sigma\circ\theta\circ\gamma_{\theta}^{-1}([-1,1])$ and $g^{j+1}\circ\sigma\circ\theta\circ\gamma_{\theta}^{1}([-1,1])$ are both much larger than $\beta$ since $\beta$ can be chosen to be much smaller than $\beta_{\varepsilon_{g}}$. Note also the image of $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is contained in the unstable manifold of $W^{u}(g^{j+1}(x))$ since the image $\sigma\circ\gamma^{\prime}\circ\theta$ is contained in the unstable manifold of $x$. This implies that $g^{j+1}(x)\in L^{u,1}(\beta)$. Thus, we get a contradiction. ∎
Thus, we have the decomposition $\Theta=\Theta_{L}\cup\Theta_{S}$, where $\theta\in\Theta_{L}$ if and only if $g^{j+1}\circ\sigma\circ\gamma^{\prime}\circ\theta$ is not $\varepsilon_{g}$-bounded. Define
$\Theta_{L}^{\pm 1}=\{\gamma^{\prime}\circ\theta\circ\gamma_{\theta}^{\pm 1}:~{}\theta\in\Theta_{L}\},$
$\Gamma_{j+1,\gamma^{\prime}}^{\prime}=\{\gamma^{\prime}\circ\theta:~{}\theta\in\Theta_{L}^{1}\cup\Theta_{L}^{-1}\cup\Theta_{S}.\},$
and
$\Gamma_{j+1}^{\prime}:=\bigcup_{\gamma^{\prime}\in\Gamma_{j}^{\prime}}\Gamma_{j+1,\gamma^{\prime}}^{\prime}.$
Thus for any $\gamma\in\Gamma_{j+1}^{\prime}$, $g^{j+1}\circ\sigma\circ\gamma$ is $\varepsilon_{g}$-bounded. From the Claim above, it is clear that
$\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{j-1},k_{j-1}^{\prime},k_{j},k_{j}^{\prime})\subset\bigcup_{\gamma\in\Gamma_{j+1}^{\prime}}\sigma\circ\gamma([-1,1]).$
By the construction, we know that the cardinality of $\Gamma_{j+1}^{\prime}$ is bounded by the quantity. ∎
Corollary 7.5.
For $n=mq$ large enough, for the family $\Gamma_{m}^{\prime}$ as in Proposition 7.4, one has that
#$\#\Gamma_{m}^{\prime}\leq 2^{m}C_{r}^{m}\exp\{n(\frac{1}{r}R(f)+\alpha/10)\}(\|Dg\|_{\rm sup}+2)^{\alpha_{1}m+1},$
Proof.
By Lemma 7.3, for $m$ large enough, one has
$\sum_{j=0}^{m-1}(k_{j}(y)-k_{j}^{\prime}(y))\leq n(\frac{1}{r}R(f)+\alpha/10)$
By the choice of $E$, i.e., $d_{m}(E)<\alpha_{1}$, one has that
#$\#\{0<i\leq m,~{}i\in E\}<\alpha_{1}m+1.$
One can thus conclude by Proposition 7.4. ∎
Now we prove Theorem 5.1 for $n=mq$. We will prove that there is a family of reparametrizations $\widetilde{\Gamma}_{mq}$ such that the following holds:
  1. 1.
    $\lim_{n\to\infty}\frac{1}{n}\log\#\widetilde{\Gamma}_{mq}\leq\frac{1}{r}R(f)+\alpha/2$;
  2. 2.
    $\bigcup_{\gamma\in\widetilde{\Gamma}_{mq}}\sigma\circ\gamma([-1,1])\supset K\cap\sigma_{*}$;
  3. 3.
    $g^{j}\circ\sigma\circ\gamma$ is strongly $\varepsilon_{g}$-bounded for any $\gamma\in\widetilde{\Gamma}_{mq}$ for any $j=0,1,\cdots,m-1$.
Now we give the proof of the above statement.
Proof.
For an admissible type $E\subset[0,m)$, for any admissible set
$\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})=\Sigma(E;k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime}),$
one gets a family $\Gamma_{j}^{\prime}=\Gamma_{j}^{\prime}(\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime}))$ of reparametrizations by applying Proposition 7.4. Thus, to take the union of all these reparametrizations, we get a family of reparametrizations $\widetilde{\Gamma}_{mq}$. By the property of $\Gamma_{j}^{\prime}$, one has that $g^{j}\circ\sigma\circ\gamma$ is strongly $\varepsilon_{g}$-bounded for any $\gamma\in\widetilde{\Gamma}_{mq}$ for any $j=0,1,\cdots,m-1$. Clearly, one has
$\bigcup_{\gamma\in\widetilde{\Gamma}_{mq}}\sigma\circ\gamma([-1,1])\supset K\cap\sigma_{*}.$
It remains to estimate the cardinality of $\widetilde{\Gamma}_{mq}$.
#$\#\widetilde{\Gamma}_{mq}\leq\#{\mathcal{S}}_{m}\times\sup_{E\in{\mathcal{S}}_{m}}\#\{(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime}):~{}\Sigma(E;k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})\neq\emptyset\}$
×$\times\sup_{E\in{\mathcal{S}}_{m};(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})}\#\Gamma_{j}^{\prime}(\Sigma(k_{0},k_{0}^{\prime},k_{1},k_{1}^{\prime},\cdots,k_{m-1},k_{m-1}^{\prime})).$
Thus, by Equation (15), Lemma 7.2, Proposition 7.4 and Corollary 7.5, one has that
#$\#\widetilde{\Gamma}_{mq}\leq{\rm e}^{\alpha n/10}\times\big{(}q\cdot\max\{\log\|Df\|_{\rm sup},~{}\log\|Df^{-1}\|_{\rm sup}\}\big{)}^{m}$
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times 2^{m}C_{r}^{m}\exp\{n(\frac{1}{r}R(f)+\alpha/10)\}(\|Dg\|_{\rm sup}+2)^{\alpha_{1}m+1}.$
By taking $\log$ and dividing $n=mq$, one has that
$\frac{1}{n}\log\#\widetilde{\Gamma}_{mq}\leq\alpha/10+\frac{1}{q}\log q+\frac{1}{q}\log\big{(}\max\{\log\|Df\|_{\rm sup},~{}\log\|Df^{-1}\|_{\rm sup}\}\big{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{q}\log(2C_{r})+\frac{1}{r}R(f)+\alpha/10+\frac{\alpha_{1}m+1}{mq}\log(\|Dg\|_{\rm sup}+2)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\frac{1}{r}R(f)+\alpha/2$
by Equation (7), Equation (4), Equation (6). ∎
Now we choose the reparametrization $\Gamma_{n}$ in Theorem 5.1 for any $n\in{\mathbb{N}}$. Write $n=mq+\ell$ with $0\leq\ell<q$. Define
$\Gamma_{n}=\widetilde{\Gamma}_{mq}.$
To prove Theorem 5.1, it remains to prove for any $n$.
$\|D(f^{j}\circ\sigma\circ\gamma)\|\leq 1,~{}~{}\forall j=0,1,\cdots,n-1.$
We have known that for any $m$, for any $\gamma\in\Gamma_{mq}$
$\|D(f^{qm}\circ\sigma\circ\gamma)\|\leq\varepsilon_{g},~{}~{}\forall j=0,1,\cdots,m-1.$
Thus, for $0\leq\ell<q$, by the choice of $\varepsilon_{g}$ (Equation 8), one has that
$\|D(f^{qm+\ell}\circ\sigma\circ\gamma)\|\leq\|Df\|_{\rm sup}^{q}\varepsilon_{g}\leq 1.$
The proof of Theorem 5.1 is thus complete. ∎
Appendix A Measurability
Given $\chi_{1}>\chi_{2}$ and $k\in{\mathbb{N}}$. We say that a point $x\in M$ has $(k,\chi_{1},\chi_{2})$-dominated Lyapunov exponents if there is a unique splitting $T_{x}M=E(x)\bigoplus F(x)$ such that $\operatorname{dim}E(x)=k$ and
  • $\lim_{n\rightarrow\pm\infty}\frac{1}{n}\log||D_{x}f^{n}|_{E(x)}||>\chi_{1}>\chi_{2}>\lim_{n\rightarrow\pm\infty}\frac{1}{n}\log||D_{x}f^{n}|_{F(x)}||.$
  • $\lim_{n\rightarrow\pm\infty}\frac{1}{n}\log\sin\angle(E(f^{n}(x)),F(f^{n}(x)))=0.$
Denote by ${\rm DOM}_{\chi_{1},\chi_{2}}^{k}$ the set of all points which have $(k,\chi_{1},\chi_{2})$-dominated Lyapunov exponents. Then ${\rm DOM}_{\chi_{1},\chi_{2}}^{k}$ is measurable and $f$-invariant.
We call that an $f$-invariant measure $\mu$ has $(k,\chi_{1},\chi_{2})$-dominated Lyapunov exponents, if there is a measurable set $\Lambda(\mu)$ with $\mu(\Lambda(\mu))=1$, $T_{\Lambda(\mu)}M:=E\bigoplus F$ such that $\operatorname{dim}E(x)=k$ and
$\lim_{n\rightarrow\pm\infty}\frac{1}{n}\log||D_{x}f^{n}(v_{E})||>\chi_{1}>\chi_{2}>\lim_{n\rightarrow\pm\infty}\frac{1}{n}\log||D_{x}f^{n}(v_{F})||$
for every $v_{E}\in E(x)\setminus\{0\},v_{F}\in F(x)\setminus\{0\}$ and every $x\in\Lambda$.
Choose $\varepsilon>0$ small enough, such that
$0<\varepsilon<\min\left\{\frac{|\chi_{1}|}{10\operatorname{dim}M},\frac{\chi_{1}-\chi_{2}}{20\operatorname{dim}M},\frac{|\chi_{2}|}{10\operatorname{dim}M}\right\}.$
By Pesin’s nonuniformly hyperbolic theory [1], there is an $f$-invariant measurable subset $\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}\subset{\rm DOM}_{\chi_{1},\chi_{2}}^{k}$, a measurable function $C:\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}\rightarrow(0,+\infty)$ and a measurable splitting $T_{\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}}=E\oplus F$ satisfying
  1. (1)
    $\mu(\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}})=1$ for any $f$-invariant measure $\mu$ which has $(k,\chi_{1},\chi_{2})$-dominated Lyapunov exponents;
  2. (2)
    $e^{-\varepsilon}<C(f(x))/C(x)<e^{\varepsilon}$ for any $x\in\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}$;
  3. (3)
    $\operatorname{dim}E(x)=k$ and $\angle(E(x),F(x))>C(x)^{-1}$ for any $x\in\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}$;
  4. (4)
    for every $x\in\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}$, for all unit vector $v_{E}\in E(x),v_{F}\in F(x)$ and for all $n\in{\mathbb{N}}$
    $||D_{x}f^{-n}v_{E}||\leq C(x)e^{-(\chi_{1}-\varepsilon)n},\ ||D_{x}f^{n}v_{F}||\leq C(x)e^{(\chi_{2}+\varepsilon)n};$
    $||D_{x}f^{n}v_{E}||\geq C(x)^{-1}e^{(\chi_{1}-\varepsilon)n},\ ||D_{x}f^{-n}v_{F}||\geq C(x)^{-1}e^{-(\chi_{2}+\varepsilon)n}.$
For $\ell>1$, we denote by
$\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell}:=\{x\in\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}:~{}C(x)\leq e^{\ell\varepsilon}\}.$
One has that $\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell}$ is closed, $\bigcup_{\ell>1}\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell}=\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}$, $f^{n}(\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell})\subset\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell+n}$ for any $n\in{\mathbb{Z}}$ and $x\mapsto E(x),x\mapsto F(x)$ depend continuously on $\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell}$ for every $n\in{\mathbb{N}}$ and $\ell>1$.
Now we assume that $\chi_{1}>0$, by the local unstable manifold theory (see [1, Section 7.1]), for any $x\in\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2}}$, the $k$-local unstable manifold $W^{u,k}_{\operatorname{loc}}(x)$ is a $C^{r}$ embedding sub-manifold, $fW^{u,k}_{\operatorname{loc}}(x)\supset W^{u}_{\operatorname{loc}}(fx)$ and $x\mapsto W^{u,k}_{\operatorname{loc}}(x)$ is continuous on $\Lambda^{k,\varepsilon}_{\chi_{1},\chi_{2},\ell}$ for any $\ell>1$. The global $k$-unstable manifold $W^{u,k}(x)$ satisfy
$W^{u,k}(x)=\bigcup_{n\geq 0}f^{n}W^{u,k}_{\operatorname{loc}}(f^{-n}x).$
For $\chi_{1}>\chi_{2}$, fix some $\varepsilon(\chi_{1},\chi_{2})$ small enough, let $\Lambda^{k}_{\chi_{1},\chi_{2}}:=\Lambda^{k,\varepsilon(\chi_{1},\chi_{2})}_{\chi_{1},\chi_{2}}$ and define $\Lambda^{k}_{\chi_{1},\chi_{2},\ell}$ similarly. Let
$\Lambda^{k}:=\bigcup\{\Lambda^{k}_{\chi_{1},\chi_{2}}:~{}\chi_{1}>\chi_{2},\ \chi_{1},\chi_{2}\in{\mathbb{Q}}\},$
and
$\Lambda^{k,u}:=\bigcup\{\Lambda^{k}_{\chi_{1},\chi_{2}}:~{}\chi_{1}>\max\{\chi_{2},0\},\ \chi_{1},\chi_{2}\in{\mathbb{Q}}\}.$
One knows that $\Lambda^{k}$ and $\Lambda^{k,u}$ are measurable. Then, for any ergodic measure $\mu$ has one-dominated positive Lyapunov exponent we have $\mu(\Lambda^{1,u})=1$.
Recall the sets $L^{u,1}(\beta)$ defined in Section 1. We now give the proof of the measurability of $L^{u,1}(\beta)$.
Proposition A.1.
The set $L^{u,1}(\beta)$ is measurable for any $\beta>0$.
Proof.
It suffice to show that for any $\chi_{1},\chi_{2}\in{\mathbb{Q}}$ and $\chi_{1}>\max\{\chi_{2},0\}$, the set $L^{u,1}(\beta)\cap\Lambda^{1}_{\chi_{1},\chi_{2}}$ is measurable.
For any $x\in\Lambda^{1}_{\chi_{1},\chi_{2}}$, we have
$W^{u,1}(x):=\bigcup_{n\geq 1}f^{n}W^{u,1}_{\rm loc}(f^{-n}x),$
and the sequence $\left\{f^{n}W^{u,1}_{\rm loc}(f^{-n}x)\right\}_{n\in{\mathbb{N}}}$ is increasing in $n$. Denote by
$~{}L^{u,1}_{n}(\beta)=$$\{x\in\Lambda^{1}_{\chi_{1},\chi_{2}}:~{}\exists W_{x}\subset f^{n}W^{u,1}_{\rm loc}(f^{-n}x),~{}\textrm{s.t.}~{}\exp_{x}^{-1}W_{x}~{}\textrm{is a $C^{1}$ graph of a map~{}}$
$~{}\varphi:~{}E(x)\to E(x)^{\perp},~{}{\rm Lip}(\varphi)\leq 1/3,~{}{\rm Domain}(\varphi)\supset E(x)(\beta)\}.$
Then, we have
$L^{u,1}(\beta)\cap\Lambda^{1}_{\chi_{1},\chi_{2}}=\bigcup_{n\geq 1,\ell\geq 1}L^{u,1}_{n}(\beta)\cap\Lambda^{1}_{\chi_{1},\chi_{2},\ell}.$ (16)
Claim.
For any $n\geq 1$ and $\ell\geq 1$, the set $L^{u,1}_{n}(\beta)\cap\Lambda^{1}_{\chi_{1},\chi_{2},\ell}$ is closed.
Proof of the Claim.
Let $(x_{m})_{m\geq 0}$ be a sequence of points in $L^{u,1}_{n}(\beta)\cap\Lambda^{1}_{\chi_{1},\chi_{2},\ell}$ and $x_{m}\rightarrow x$. Since $\Lambda^{1}_{\chi_{1},\chi_{2},\ell}$ is closed, one has $x\in\Lambda^{1}_{\chi_{1},\chi_{2},\ell}$. Note that $f^{-n}(x_{m})\in\Lambda^{1}_{\chi_{1},\chi_{2},\ell+n}$ and $f^{-n}(x_{m})\rightarrow f^{-n}(x)$, one has $W^{u,1}_{\rm loc}(f^{-n}(x_{m}))$ converge to $W^{u,1}_{\rm loc}(f^{-n}(x))$ as $m\rightarrow+\infty$ in the $C^{1}$ topology. So, we also have $f^{n}W^{u,1}_{\rm loc}(f^{-n}(x_{m}))$ converge to $f^{n}W^{u,1}_{\rm loc}(f^{-n}(x))$ as $k\rightarrow+\infty$ in the $C^{1}$ topology.
Note that $E(x_{m})$ also converge to $E(x)$. By $\{x_{m}\}_{m\in{\mathbb{N}}}\subset L^{u}_{n}(\beta)$, one can easily to prove that there exists a map $\varphi:~{}E(x)\to E(x)^{\perp}$ with ${\rm Domain}(\phi)\supset E(x)(\beta)$ and ${\rm Lip}(\varphi)\leq 1/3$, such that $\exp_{x}{\rm graph}(\varphi)\subset f^{n}W^{u,1}_{\rm loc}(f^{-n}(x))$. Thus, $x\in L^{u,1}_{n}(\beta)$ and so $L^{u,1}_{n}(\beta)\cap\Lambda^{1}_{\chi_{1},\chi_{2},\ell}$ is closed. ∎
The proof of the proposition follows from the claim and Equation (16) immediately. ∎